# Verification methods for international AI agreements

### v1

@@ Introduction
If the total amount of energy reaching a data center can be measured with reasonable accuracy, it should be possible to convert the energy estimate into a reasonable approximation of the number of FLOPs completed by that facility (Desislavov, Martínez-Plumed, and Hernández-Orallo 2021). Using the FLOP/s, we could ascertain whether the facility is at an unauthorized size. Although much harder, with a sufficient degree of sensitivity, it may also be possible to use energy monitoring to determine whether a data center is hosting an unauthorized training run. For example, there could be a shift in properties of the power consumption as the data center transitions to a more constant GPU utilization during training when compared to a variable number of inference requests.
Precedent. Economists use energy monitoring to verify economic data. For example, Thomas Rawski used discrepancies between reported GDP growth and energy consumption to suggest exaggerated growth figures in China (Owyang and Shell 2017). This principle could be applied to detect unauthorized data centers, given the direct relationship between energy consumption and FLOPS.
@@
An international inspection team could visit an AI development facility to ensure that developers are running authorized code, ensure that developers are properly implementing model evaluations and safeguards, assess safety culture, and assess security concerns. Inspections could involve various components, such as reviewing code (Casper et al. 2024), assessing compliance with commitments from safety cases555For example, whether mandated interpretability techniques are implemented (Räuker et al. 2023; Bereska and Gavves 2024), or evaluations (Shevlane et al. 2023; Phuong et al. 2024; Anderljung et al. 2023)., and conducting semi-structured interviews with key personnel to solicit security-relevant concerns (see Wasil et al. (2024a)). Inspections could uncover the usage of unauthorized or unlicensed AI algorithms.

### v2

@@ Introduction
It is important to note that incentivization alone may not be sufficient to ensure the effectiveness of whistleblower schemes, given that determined adversaries might attempt to physically or digitally block employees from contacting a verifying authority.
One possible solution to this limitation is to implement regular in-person communication with employees, such as through semi-structured interviews (Wasil et al. 2024a).555We thank Mauricio Baker for this suggestion, which is also included in his upcoming report.
@@
If the total amount of energy reaching a data center can be measured with reasonable accuracy, it should be possible to convert the energy estimate into a reasonable approximation of the number of FLOPs completed by that facility (Desislavov, Martínez-Plumed, and Hernández-Orallo 2021). Using the FLOP/s, we could ascertain whether the facility is at an unauthorized size. However, these coarse-grained measurements may only be capable of detecting large-scale violations, and further research is needed to understand how such measurements can be more accurately translated into relevant units like FLOPs. More precise methods of energy monitoring may be necessary for detecting smaller-scale violations or unauthorized training runs.
Precedent. Economists use energy monitoring to verify economic data. For example, economists have used discrepancies between reported GDP growth and energy consumption to suggest exaggerated growth figures in China (Owyang and Shell 2017). This principle could be applied to detect unauthorized data centers, given the direct relationship between energy consumption and FLOPS.
@@
An international inspection team could visit an AI development facility to ensure that developers are running authorized code, properly implementing model evaluations and safeguards, and assess safety culture and security concerns. Inspections could involve various components, such as reviewing code (Casper et al. 2024), assessing compliance with commitments from safety cases666For example, whether mandated interpretability techniques are implemented (Räuker et al. 2023; Bereska and Gavves 2024), or evaluations (Shevlane et al. 2023; Phuong et al. 2024; Anderljung et al. 2023)., and conducting semi-structured interviews with key personnel to solicit security-relevant concerns (see Wasil et al. (2024a)). Inspections could uncover the usage of unauthorized or unlicensed AI algorithms. A number of privacy-preserving technologies in development could facilitate such inspections without being overly intrusive. .777We thank Mauricio Baker for this suggestion, which is also included in his upcoming report.